Three vectors ,x,y, and z satisfy x+y+z=0 . Calculate the value of


 x. y+y.z+x.z  if lxl =2 , lyl=3 and lzl=4

Grade:

2 Answers

Abhishek V
13 years ago

14.5

vikas askiitian expert
13 years ago

x + y + z = 0

squaring both sides

 

x2 + y2 + z2 + 2(x.y+y.z+z.x) = 0

 

x.y + y.z + z.x = (x2+y2+z2)/2               ..............1

now plugging values of x,y z we get

 

x.y + y.z + z.x = 29/2

 

approve if u like my ans